OpenStudy (anonymous):

*MEDAL* Solve for x and y. -3x - 2y = -12

OpenStudy (anonymous):

For y, You isolate the y variable. -3x-2y=-12 -3x+12=2y y = -3/2x+6

OpenStudy (anonymous):

For x, isolate the x variable -3x=-12+2y x=4-2/3y

OpenStudy (anonymous):

Original Question:

OpenStudy (anonymous):

Choices: (2, 1) no solutions (-2, -1) infinitely many solutions

OpenStudy (anonymous):

@campbell_st Could you help with this one? Once I get this last question I can finally go to bed :/ (Doing work since 8:30pm to now, 8:43pm)

OpenStudy (anonymous):

Nvm heading to bed none-the-less :) thanks anyways!